# A Systematic Review of Aspect-based Sentiment Analysis (ABSA): Domains, Methods, and Trends thanks: Preprint manuscript

### v4

@@ 5 Discussion
This review was motivated by the literature gap in capturing trends in ABSA research to answer higher-level questions beyond technical details, and the concern that the domain-dependent nature could predispose ABSA research to systemic hindrance from a combination of resource-reliant approaches and skewed resource domain distribution. By systematically reviewing the 519 in-scope primary studies published between 2008 and 2022, our quantitative analysis results identified trends in ABSA solution approaches, confirmed the above-mentioned concern, and provided detailed insights into the relevant issues. In this section, we examine the primary findings, share ideas for future research, and reflect on the limitations of this review.
@@ 5.1.3 The Reliance on Labelled-Datasets
As demonstrated in Section 2.1.2, the domain can directly affect whether a chunk of text is considered an aspect or the relevant sentiment term, and plays a crucial role in contextual inferences such as implicit aspect extraction and multi-aspect multi-sentiment pairing. The domain knowledge reflected via ABSA labelled datasets can further shape the linguistic rules, lexicons, and knowledge graphs for non-machine-learning approaches; and define the underpinning feature space, representations, and acquired relationships and inferences for trained machine-learning models. When applying a solution built on datasets from a domain that is very remote from or much narrower than the intended application domain, it is predictable that the solution performance would be capped at subpar and even fail at more context-heavy tasks [92, 21, 11, 88, 93, 82, 19]. Thus, domain transfer is crucially necessary for balancing the uneven ABSA research and resource distributions across domains. However, our finding that 17 out of the 20 reviewed cross-domain or domain-agnostic ABSA studies solely used datasets from the “product/service review” domain raised questions about these approaches’ generalisability and robustness in other domains, as well as whether such dataset choices became another reinforcement of concentrating research effort and benchmarks within this one dominant domain.
@@
Moreover, RNN-based solutions dominate the DL approaches (55.91%, N=175), mainly with the RNN-attention combination (26.52%, N=83) and RNN-only (9.90%, N=31) models. RNN and its variants such as Long short-term memory (LSTM) and Gated recurrent units (GRUs) are known for their limitations in capturing long-distance relations due to their sequential nature and the subsequent memory constraints [43, 20]. Although the addition of the attention mechanism enhances the model’s focus on more important features such as aspect terms [43, 20], traditional attention weights calculation struggles with multi-word aspects or multi-aspect sentences [20, 103]. In addition, whilst 16.77% (N=53) of the fully-supervised studies introduced syntactical features to their DL solutions, additional features also increased the input size. According to [104], sequential models, even the state-of-the-art large language models, showed impaired performance as the input grew longer and could not always benefit from additional features.
@@ 5.2 Ideas for Future Research
Overall, by adopting a “systematic perspective, i.e., model, data, and training” [94,  p.28] combined with a quantitative approach, we identified high-level trends unveiling the development and direction of ABSA research, and found clear evidence of large-scale issues that affect the majority of the existing ABSA research. The skewed domain distributions of resources and benchmarks could also restrict the choice of new studies. On the other hand, this evidence also highlights areas that need more attention and exploration, such as ABSA solutions and resource development for the less-studied domains (e.g. education and public health), low-resource and/or data-agnostic ABSA, domain adaptation, alternative training schemes such as adversarial (e.g. [94, 105]) and reinforcement learning (e.g. [106, 107]), and more effective feature and knowledge injection.
In addition, our results also revealed emerging trends and new ideas. The relatively recent growth of end-to-end models and composite ABSA subtasks provide opportunities for further exploration and evaluation. The fact that hybrid approaches with non-machine-learning techniques and non-textual features remain steady forces in the field after nearly three decades suggests valuable characteristics that are worth re-examining under the light of new paradigms and techniques. Moreover, this review was conducted before the mass popularity of generative language models, such as ChatGPT [108]. It is worth exploring how ABSA research could ethically leverage the capabilities of pre-trained generative models, potentially for both ABSA solution and resource creation.
@@ 5.3 Limitations
We acknowledge the following limitations of this review: First, our sample scope is by no means exhaustive, as it only includes primary studies from four peer-reviewed digital databases and only those published in the English language. Although this can be representative of a core proportion of ABSA research, it does not generalise beyond this without assumptions. Second, no search string is perfect. Our database search syntax and auto-screening keywords represent our best effort in capturing ABSA primary studies, but may have missed some relevant ones, especially with the artificial “total pages <<< 3” and “total keyword (except SA, OM) outside Reference <<< 5” exclusion criteria. Third, we may have missed datasets, paradigms, and approaches that are not clearly described in the primary studies, and our categorisation of them is also subject to the limitations of our knowledge and decisions. Fourth, we did not compare solution performance across studies given the focus of this review and the scale of our sample.
@@ 6 Conclusion
Appendix A Appendix - Full SLR Methodology
@@
A.1 Research identification
To obtain the files for review, we conducted database searches between 24-25 October 2022, when we manually queried and exported a total of 4191 research papers’ PDF and BibTeX (or the equivalent) files via the web interfaces of four databases. Table A.1 details the search string, search criteria, and the PDF files exported from each database.
Table A.1: Digital databases and search details used for this Systematic Literature Review (SLR)
@@ A.1 Research identification
A.2 Selection of primary studies
@@
To implement the inclusion and exclusion criteria, we first applied PDF mining to automatically exclude files that meet the exclusion criteria, and then refined the selection with manual screening under the exclusion and inclusion criteria. Both of these processes are detailed below.
The automatic screening consists of a pipeline with two Python packages: Pandas [110] and PyMuPDF [^7] https://pypi.org/project/PyMuPDF/. We first used Pandas to extract into a dataframe (i.e. table) all exported papers’ file locations and key BibTex or equivalent information including title, year, page number, DOI, and ISBN. Next, we used PyMuPDF to iterate through each PDF file and add to the dataframe multiple data fields: whether the file was successfully decoded [^8] https://pdfminersix.readthedocs.io/en/latest/faq.html for text extraction (if marked unsuccessful, the file was marked for manual screening), the occurrence count of each Regex keyword pattern listed below, and whether each keyword occurs after the section headings that fit into Regex patterns that represent variations of “references” and “bibliography” (referred to as “non-target sections” below). We then marked the files for exclusion by evaluating the eight criteria listed under “Auto-excluded” in Table A.2 against the information recorded in the dataframe. Each of the auto-exclusion results from Steps 1-4 and 7 in Table A.2 were manually checked, and those under Steps 5, 6, and 8 were spot-checked. These steps excluded 3277 out of the 4194 exported files.
@@
For the 914 files filtered through the auto-exclusion process, we manually screened them individually according to the inclusion and exclusion criteria. As shown in the second half of Table A.2, this final screening step refined the review scope to 519 papers.
@@
Table A.2: The automatic and manual screening processes
@@
A.3 Data extraction and synthesis
@@
Appendix B Appendix - Additional Results
Figure B.1: Number of Included Studies by Publication Year and Type (N=519)

### v5

@@ 5 Discussion
This review was motivated by the literature gap in capturing trends in ABSA research to answer higher-level questions beyond technical details, and the concern that the domain-dependent nature could predispose ABSA research to systemic hindrance from a combination of resource-reliant approaches and skewed resource domain distribution. By systematically reviewing the two waves of 727 in-scope primary studies published between 2008 and 2024, our quantitative analysis results identified trends in ABSA solution approaches, confirmed the above-mentioned concern, and provided detailed insights into the relevant issues. In this section, we examine the primary findings, share ideas for future research, and reflect on the limitations of this review.
@@ 5.1.3 The Reliance on Labelled-Datasets
As demonstrated in Section 2.3, the domain can directly affect whether a chunk of text is considered an aspect or the relevant sentiment term, and plays a crucial role in contextual inferences such as implicit aspect extraction and multi-aspect multi-sentiment pairing. The domain knowledge reflected via ABSA labelled datasets can further shape the linguistic rules, lexicons, and knowledge graphs for non-machine-learning approaches; and define the underpinning feature space, representations, and acquired relationships and inferences for trained machine-learning models. When applying a solution built on datasets from a domain that is very remote from or much narrower than the intended application domain, it is predictable that the solution performance would be capped at subpar and even fail at more context-heavy tasks [52, 21, 11, 48, 53, 38, 19]. Thus, domain transfer is crucially necessary for balancing the uneven ABSA research and resource distributions across domains. However, our finding that 17 out of the 20 reviewed cross-domain or domain-agnostic ABSA studies solely used datasets from the “product/service review” domain raised questions about these approaches’ generalisability and robustness in other domains, as well as whether such dataset choices became another reinforcement of concentrating research effort and benchmarks within this one dominant domain.
@@
Moreover, RNN-based solutions dominate the DL approaches (55.91%, N=175), mainly with the RNN-attention combination (26.52%, N=83) and RNN-only (9.90%, N=31) models. RNN and its variants such as LSTM, BiLSTM, and GRU are known for their limitations in capturing long-distance relations due to their sequential nature and the subsequent memory constraints [74, 20]. Although the addition of the attention mechanism enhances the model’s focus on more important features such as aspect terms [74, 20], traditional attention weights calculation struggles with multi-word aspects or multi-aspect sentences [20, 121]. In addition, whilst 16.77% (N=53) of the fully-supervised studies introduced syntactical features to their DL solutions, additional features also increased the input size. According to [122], sequential models, even the state-of-the-art LLMs, showed impaired performance as the input grew longer and could not always benefit from additional features.
5.1.4 The Potential of Generative LLMs and Foundation Models
Lastly, the Phase-2 targeted review highlights the ABSA community’s caution toward the direct adoption of generative foundation models, with only five out of 208 recent studies testing the ICL approach and most yielding subpar results compared to other methods. However, most of these studies only tested zero-shot instructions with simple model settings. It is worth further exploring the potential of foundational models and ICL in ABSA by focusing more on instruction and example engineering, model parameter optimisation, and task re-formulation [88].
On the other hand, the fine-tuning of smaller generative LLMs has seen increasing adoption through the “ABSA as Seq2Seq text generation” approach, demonstrating promising task performance. Although this generative approach can incorporate data augmentation and self-training to reduce reliance on labelled datasets, the cost of fine-tuning, the need for labelled base data, and the domain-transfer problem remain significant challenges [73]. In this context, the task adaptability and multi-domain pre-trained knowledge of foundation models could provide potential solutions.
As Zhang et al. (2022) [73] noted, progress in applying pre-trained LLMs and foundation models to ABSA could be impeded by dataset resources constraints. To match the parameter size of these models, more diverse, complex, and larger datasets are required for effective fine-tuning or comprehensive testing. In low-resource domains where dataset resources are already limited, this requirement could further complicate the adoption of these technologies [54].
@@ 5.2 Ideas for Future Research
Overall, by adopting a “systematic perspective, i.e., model, data, and training” [55,  p.28] combined with a quantitative approach, we identified high-level trends unveiling the development and direction of ABSA research, and found clear evidence of large-scale issues that affect the majority of the existing ABSA research. The skewed domain distributions of resources and benchmarks could also restrict the choice of new studies. On the other hand, this evidence also highlights areas that need more attention and exploration, including: ABSA solutions and resource development for the less-studied domains (e.g. education and public health), low-resource and/or data-agnostic ABSA, domain adaptation, alternative training schemes such as adversarial (e.g. [55, 123]) and reinforcement learning (e.g. [124, 125]), and more effective feature and knowledge injection. Future research could contribute to addressing these issues by focusing on ethically producing and sharing more diverse and challenging datasets in minority domains such as education and public health, improving data synthesis and augmentation techniques, exploring methods that are less data-dependent and resource-intensive, and leveraging the rapid advancements in pre-trained LLMs and foundation models.
In addition, our results also revealed emerging trends and new ideas. The relatively recent growth of end-to-end models and composite ABSA subtasks provide opportunities for further exploration and evaluation. The fact that hybrid approaches with non-machine-learning techniques and non-textual features remain steady forces in the field after nearly three decades suggests valuable characteristics that are worth re-examining under the light of new paradigms and techniques. Moreover, the small number of Phase-2 samples using ICL and fine-tuning generative LLM approaches may suggest that we have only captured early adopters. More thorough exploration of these approaches and continued tracking of their development alongside other methods are necessary to understand how the ABSA community can leverage the resources and capabilities embedded within LLMs and foundation models.
@@ 5.3 Limitations
We acknowledge the following limitations of this review: First, our sample scope is by no means exhaustive, as it only includes primary studies from four peer-reviewed digital databases and only those published in the English language. Although this can be representative of a core proportion of ABSA research, it does not generalise beyond this without assumptions. The “peer-reviewed” criteria also meant that we overlooked preprint servers such as arXiv.org that more closely track the latest development of ML and NLP research. Second, no search string is perfect. Our database search syntax and auto-screening keywords represent our best effort in capturing ABSA primary studies, but may have missed some relevant ones, especially with the artificial “total pages <<< 3” and “total keyword (except SA, OM) outside Reference <<< 5” exclusion criteria. Moreover, our search completeness might have been affected by the performance of the database search engines. This is evidenced by the significant number of extracted search results that were entirely irrelevant to the search keywords, as well as our abandonment of the 2024 SpringerLink search due to interface issues. Enhancements in digital database search capabilities could significantly improve the effectiveness and reliability of future literature review studies, particularly SLRs. Third, we may have missed datasets, paradigms, and approaches that are not clearly described in the primary studies, and our categorisation of them is also subject to the limitations of our knowledge and decisions. Future review studies could consider a more innovative approach to enhance analytical precision and efficiency, such as applying ABSA and text summarisation alongside the screening and reviewing process. Fourth, we did not compare solution performance across studies due to the review focus, sample size, and the variability in experimental settings across studies. Evaluating the effectiveness of comparable methods and the suitability of evaluation metrics would enhance our findings and offer more valuable insights.
@@ 6 Conclusion
Appendix A Aspect-Based Sentiment Analysis (ABSA)
A.1 Definition and examples
The following examples illustrate the ABSA terminologies:
As demonstrated above, the fine granularity makes ABSA more targetable and informative than document- or sentence-level SA. Thus, ABSA can precede downstream applications such as attribute weighting in overall review ratings (e.g. [127]), aspect-based opinion summarisation (e.g. [128, 129, 130], and automated personalised recommendation systems (e.g. [131, 97]).
A.2 ABSA Subtasks
With the rise of multi-task learning and deep learning [143], an increasing number of studies explore ABSA under an End-to-end (E2E) framework that performs multiple fundamental ABSA subtasks in one model to better capture the inter-task relations [37], and some combine them into a single composite task [29, 35, 38]. These composite tasks are most commonly formulated as a sequence- or span-based tagging problem [29, 35, 36]. The most common composite tasks are: Aspect-Opinion Pair Extraction (AOPE), which directly outputs {aspect, opinion} pairs from text input [36, 39, 40] such as “⟨menu, great⟩” from Example 1; Aspect-Polarity Co-Extraction (APCE) [29, 41], which outputs {aspect, sentiment polarity} pairs such as “⟨menu, positive⟩”; Aspect-Sentiment Triplet Extraction (ASTE) [29, 35, 42, 31], which outputs {aspect, opinion, sentiment category} triplets, such as “⟨menu, great, positive⟩”; and Aspect-Sentiment Quadruplet Extraction/Prediction (ASQE/ASQP) [43, 44, 45, 46] that outputs {aspect, opinion, aspect category, sentiment category} quadruplets, such as “⟨menu, great, general, positive⟩”.
A.3 Other ABSA Reviews
As this review focuses on trends instead of detailed solutions and methodologies, we refer interested readers to existing review papers that provide comprehensive and in-depth summaries of common ABSA subtask solutions and approaches, for example:
  • Explicit and implicit AE:  Rana and Cheah (2016) [56], Ganganwar and Rajalakshmi (2019) [68], Soni and Rambola (2022) [67], Maitama et al. (2020) [24]
  • Deep learning (DL) methods for ABSA:  Do et al. (2019) [57], Liu et al. (2020) [20], Wang et al. (2021) [70], Chen and Fnu (2022) [71], Zhang et al. (2022) [73], Mughal et al. (2024) [72].
    Specifically:
      DL methods for ASC:  Zhou et al. (2019) [69], Satyarthi and Sharma (2023) [54]
      E2E ABSA, composite tasks, and pre-trained Large Language Models (LLMs) in ABSA:  Zhang et al. (2022) [73] provided a comprehensive review and shared extensive reading lists and dataset resource links via https://github.com/IsakZhang/ABSA-Survey. Mughal et al. (2024) [72] introduced common benchmark datasets, including more challenging ones for composite ABSA tasks. They also reviewed and tested the ABSA task performance of representative RNN-based models and pre-trained LLMs.
  • Multimodal ABSA:  Zhao et al. (2024) [144]
Appendix B Full SLR Methodology
@@
B.1 Research identification
To obtain the files for review, we conducted database searches between 24-25 October 2022, when we manually queried and exported a total of 4191 research papers’ PDF and BibTeX (or the equivalent) files via the web interfaces of four databases. Table B.1 details the search string, search criteria, and the PDF files exported from each database.
Table B.1: Digital databases and search details used for this Systematic Literature Review (SLR)
@@ B.1 Research identification
B.2 Selection of primary studies
@@
To implement the inclusion and exclusion criteria, we first applied PDF mining to automatically exclude files that meet the exclusion criteria, and then refined the selection with manual screening under the exclusion and inclusion criteria. Both of these processes are detailed below. Our PDF mining for automatic review screening code is also available at https://doi.org/10.5281/zenodo.12872948.
The automatic screening consists of a pipeline with two Python packages: Pandas [146] and PyMuPDF [^9] https://pypi.org/project/PyMuPDF/. We first used Pandas to extract into a dataframe (i.e. table) all exported papers’ file locations and key BibTex or equivalent information including title, year, page number, DOI, and ISBN. Next, we used PyMuPDF to iterate through each PDF file and add to the dataframe multiple data fields: whether the file was successfully decoded [^10] https://pdfminersix.readthedocs.io/en/latest/faq.html for text extraction (if marked unsuccessful, the file was marked for manual screening), the occurrence count of each Regex keyword pattern listed below, and whether each keyword occurs after the section headings that fit into Regex patterns that represent variations of “references” and “bibliography” (referred to as “non-target sections” below). We then marked the files for exclusion by evaluating the eight criteria listed under “Auto-excluded” in Table B.2 against the information recorded in the dataframe. Each of the auto-exclusion results from Steps 1-4 and 7 in Table B.2 were manually checked, and those under Steps 5, 6, and 8 were spot-checked. These steps excluded 3277 out of the 4194 exported files.
@@
For the 914 files filtered through the auto-exclusion process, we manually screened them individually according to the inclusion and exclusion criteria. As shown in the second half of Table B.2, this final screening step refined the review scope to 519 papers.
@@
Table B.2: The automatic and manual screening processes
@@
B.3 Data extraction and synthesis
@@
Appendix C Additional Results
Figure C.1: Number of Included Studies by Publication Year and Type (N=519)
@@ Appendix C Additional Results
Figure C.2: Number of Included Studies with the Top 5 Dataset Languages by Publication Year
Table C.1: Number of Studies by ABSA Subtask Combinations (N=519)
Subtask Combination Count of Studies % of Studies
AE, ASC 168 32.37%
ASC 160 30.83%
AE 79 15.22%
ACD, ASC 16 3.08%
ASTE 15 2.89%
AE, ACD, ASC 13 2.50%
AE, OE, ASC 9 1.73%
AE, ACD 9 1.73%
AOPE 9 1.73%
ACD 7 1.35%
AE, OE 7 1.35%
AE, OE, ACD 4 0.77%
AE, OE, ACD, ASC 4 0.77%
ASQE 2 0.39%
AOPE, ASC 2 0.39%
OE 1 0.19%
OTE (Opinion Target Extraction), ASC 1 0.19%
Aspect-based embedding, ACD, ASC 1 0.19%
ASC, OE 1 0.19%
Aspect and synthetic sample discrimination 1 0.19%
AE, OE, ASC, AOPE, ASTE 1 0.19%
AOPE, ASC, ACD 1 0.19%
AOPE, ACD 1 0.19%
AE, review-level SA 1 0.19%
ACD, AE, ASC 1 0.19%
AE, ASC, Aspect-based sentence segmentation 1 0.19%
AE, ASC, Aspect-based embedding 1 0.19%
AE, ASC, ACD 1 0.19%
AE, ACD, OE 1 0.19%
Data augmentation 1 0.19%
TOTAL 519 100.00%
  • 1
    This table corresponds to Figure  5(a)
Table C.2: Number of Studies by Individual ABSA Subtasks (N=805)
Individual Subtask Count of Studies % of Studies
ASC 381 47.33%
AE 300 37.27%
ACD 59 7.33%
OE 28 3.48%
ASTE 16 1.99%
AOPE 14 1.74%
ASQE 2 0.25%
Aspect and synthetic sample discrimination 1 0.12%
OTE (Opinion Target Extraction) 1 0.12%
Aspect-based sentence segmentation 1 0.12%
Data augmentation 1 0.12%
Review-level SA 1 0.12%
TOTAL 805 100%
  • 1
    This table corresponds to Figure  5(b)
Table C.3: Number of Studies by Deep Learning (DL) Approaches (N=313)
DL Approach Count of Studies % of Studies
RNN, Attention 83 26.52%
Attention 60 19.17%
RNN 31 9.90%
RNN, CNN, Attention 25 7.99%
CNN 20 6.39%
RNN, CNN 19 6.07%
GNN/GCN, Attention 17 5.43%
CNN, Attention 17 5.43%
RNN, GNN/GCN, Attention 17 5.43%
Other (<5% each) 24 7.67%
TOTAL 313 100.00%
  • 1
    This table corresponds to Figure  7(a)
Table C.4: Number of Studies by Traditional Machine Learning Approaches (N=283)
Traditional ML Count of Studies % of Studies
Support Vector Machine (SVM) 57 20.14%
Conditional Random Field (CRF) 41 14.49%
Latent Dirichlet allocation (LDA) 36 12.72%
Naïve Bayes (NB) 32 11.31%
Random Forest (RF) 24 8.48%
Decision Tree (DT) 18 6.36%
Logistic Regression (LR) 15 5.30%
K-Nearest Neighbors (KNN) 12 4.24%
Multinomial Naïve Bayes (MNB) 10 3.53%
K-means 5 1.77%
Boosting 5 1.77%
Other (N<5) 28 9.89%
TOTAL 283 100.00%
  • 1
    This table corresponds to Figure  7(b)
Table C.5: Number of Studies Per Dataset (N=1179, with 519 studies and 218 datasets)
Datasets Count of Studies % of Studies
SemEval 2014 Restaurant 211 17.90%
SemEval 2014 Laptop 189 16.03%
SemEval 2016 Restaurant 118 10.01%
SemEval 2015 Restaurant 106 8.99%
Twitter (Dong et al., 2014) 55 4.66%
Amazon customer review datasets (Hu & Liu, 2004) 33 2.80%
Amazon product review (original) 23 1.95%
Product review (original) 22 1.87%
Twitter (original) 21 1.78%
SemEval 2015 Laptop 20 1.70%
Yelp Dataset Challenge Reviews 16 1.36%
SemEval 2016 Laptop 14 1.19%
TripAdvisor Hotel review (original) 11 0.93%
Hotel review (original) 10 0.85%
TripAdvisor Restaurant review (original) 9 0.76%
Movie review (original) 8 0.68%
MAMS Multi-Aspect Multi-Sentiment dataset (Jiang et al., 2019) 8 0.68%
SemEval 2016 Hotel 7 0.59%
Twitter (Mitchell et al., 2013) 7 0.59%
Amazon product review (McAuley, 2016) 7 0.59%
VLSP 2018 Restaurant review 6 0.51%
Restaurant review (Ganu et al., 2009) 6 0.51%
Student feedback (original) 6 0.51%
VLSP 2018 Hotel review 6 0.51%
TripAdvisor hotel review (Want et al., 2010) 5 0.42%
Chinese product review (Peng et al., 2018) 5 0.42%
Yelp restaurant review (original) 5 0.42%
SentiHood (Saeidi et al., 2016) 4 0.34%
TripAdvisor hotel review (Wang et al., 2010) 4 0.34%
Game review (original) 3 0.25%
Sanders Twitter Corpus (STC) (Sanders, 2013) 3 0.25%
TripAdvisor Tourist review (original) 3 0.25%
Coursera course review (original) 3 0.25%
Online drug review (original) 3 0.25%
Stanford Twitter Sentiment (STS) 3 0.25%
Restaurant review (original) 3 0.25%
Chinese Restaurant review (original) 3 0.25%
Financial Tweets and News Headlines dataset (FiQA 2018 ) 2 0.17%
SentiRuEval-2015 2 0.17%
Amazon laptop reviews (He & McAuley, 2016) 2 0.17%
Hindi ABSA dataset (Akhtar et al., 2016) 2 0.17%
Sentiment140 dataset (Go et al.) 2 0.17%
Hotel review (Kaggle) 2 0.17%
Service review (Toprak etc. 2010) 2 0.17%
Indonesian Product review (original) 2 0.17%
Indonesian Tweets - Political topic (original) 2 0.17%
Steam game review (original) 2 0.17%
Stanford Sentiment Treebank review data (Socher, 2013) 2 0.17%
Kaggle movie review 2 0.17%
Zomato restaurant review (original) 2 0.17%
Turkish Product review (original) 2 0.17%
SemEval 2015 Hotel 2 0.17%
Weibo comments (original) 2 0.17%
Amazon 50-domain Product review (Chen & Liu, 2014) 2 0.17%
Bangla ABSA Cricket, Restaurant dataset (Rahman & Dey, 2018) 2 0.17%
YouTube comments (original) 2 0.17%
Beer review (McAuley et al., 2012) 2 0.17%
CCF BDCI 2018 Chinese auto review dataset 2 0.17%
ReLi Portuguese book reviews (Freitas, et al., 2012) 2 0.17%
Twitter product review (original) 2 0.17%
SemEval 2016 Tweets 2 0.17%
Product review dataset (Liu et al., 2015) 2 0.17%
Product review dataset (Cruz-Garcia et al., 2014) 2 0.17%
Chinese Product review (original) 2 0.17%
ICLR Open Reviews dataset 2 0.17%
Amazon Electronics review data (Jo, 2011) 2 0.17%
Amazon product review (Wang et al., 2011) 2 0.17%
SemEval 2015 (unspecified) 2 0.17%
Product & service review (Kaggle) 2 0.17%
Othe datasets (N=1 each) 149 12.64%
TOTAL 1179 100.00%